Consider robin’s eggs; the mean weight of a collection of robin eggs is 0.72 ounce and the standard deviation is 0.12 ounce. Now consider elephants; the men weight of elephants in the zoo is 6.42 tons, with a standard deviation of 1.07 tons. The units of measurement are different and there is a great deal of difference between the weight of an elephant and that of a robin’s egg. Yet, the coefficient of variation is about the same for both?
Answer #1

a) let X : robin's egg weight has mean x = 0.72 ounce and standard deviation x = 0.12 ounce

Y : elephant weight has mean x = 6.42 tons and standard deviation x = 1.07 tons

The coefficient of variation for robin's egg weight is

C . V for robin =( x /x )*100

C.V for robin = (0.12/0.72)*100

C.V for robin = 16.67%

Coefficient of variation for elephants weight

C.V for elephant = (1.07/6.42)*100

C.V for elephant = 16.67%

Yes ,coefficient of variation is same for both.